OpenStudy (anonymous):

Find the midpoint of segment P Q if P(-8, -8) and Q(4, 4)? Answer (2, -2) (-2, -2) (2, 2) (1, 2)

hartnn (hartnn):

The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

(2,2) ?

hartnn (hartnn):

what is (-8+4)/2 =... ? its not 2....

hartnn (hartnn):

whats -8+4 =... ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

2

OpenStudy (anonymous):

-6

hartnn (hartnn):

guessing ? -8+4 = -4! -4 /2 = -2 so, (-2,-2)

OpenStudy (primeralph):

This is nice.